# Towards Terminology Management Automation for Arabic

### v1

@@ Abstract
This paper presents a method and supporting tools for automation of terminology management for Arabic. The tools extract lists of parallel terminology matching terms in foreign languages to their Arabic counterparts from field specific texts. This has significant implications as it can be used to improve consistent translation and use of terms in specialized Arabic academic books, and provides automated aid for enhancing cross-lingual text processing. This automation of terminology management aims to reduce processing time, and ensure use of consistent and correct terminology.
The extraction takes advantage of naturally occurring term translations. It considers several candidate phrases of varying lengths that co-occur next to the foreign terms.
Then it computes several similarity metrics, including lexicographic, phonetic, morphological, and semantic ones to decide the problem. We experiment with heuristic, machine learning, and ML with post processing approaches. This paper reports on a novel curated dataset for the task, an existing expert reviewed industry parallel corpora, and on the performance of the three approaches. The best approach achieved $94.9\%$ precision and $92.4\%$ recall.
Towards Terminology Management Automation for Arabic
@@ 1 Introduction
Translation of academic literature is essential to knowledge transfer. Linguistic and structural complexities exist. These affect consistency and clarity of translated material and may hinder knowledge accessibility. Terminology management concerns the complexities involved in (i) specifying the term that should be used upon the introduction of a new concept to a language, and (ii) and enforcing the use, hence (iii) ensuring consistency and correctness.
The first and most important step is to build TermBase, a large database of parallel terms. Authors can utilize TermBase to enhance their writing processes when faced with a new concept; be it a foreign term to be translated to Arabic.
@@
Recent advances in natural language processing and computational linguistics have paved the way for automated systems that assist in cross-lingual text processing. However, existing systems, including LLMs, often struggle with the nuances of terminology in cross-lingual specific expert fields, where precision is paramount. This paper introduces a novel methodology and a supporting framework to extract terminology mappings from existing Arabic books, a rich source of terms when building a TermBase.
@@
When a foreign term occurs in Arabic books, usually between parentheses, it is preceded by its Arabic translation. Two problems arise. The exact scope of the Arabic term is ambiguous, and often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones. Editors and reviewers, who are usually expert researchers, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
@@
Our work helps automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
@@
The problem is formulated as follows: given a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$, denote candidate terms as the set of strings $s_{i}=w_{1}\ldots w_{i}$ ending before term $f$, $1\leq i\leq n$ as a candidate term. The goal is to identify the target candidate term $s_{t}$ best matching $f$. Our tool extracts linguistic features for each candidate match $(s_{i},f)$. The tool interpolates the features and aggregates their values into a resulting score that ranks the candidate terms.
@@
By addressing the challenges of terminology mapping, this research moves us closer to a more integrated community where language barriers are significantly reduced, facilitating a greater exchange of knowledge and ideas across languages and cultures.
@@
We make verified parts of the constructed dataset available online for the research community. To the best of our knowledge, this approach to TermBase extraction for Arabic is novel, and no comparable datasets currently exist.
@@ 1.1 Motivating example
Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme with different translations.
@@
Candidates for Context 1:
@@
  • We create a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
  • We evaluate and compare three approaches: heuristic, machine learning, and machine learning with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges. The results are impressive and convincing for deployment. The final model runs locally, requires no LLM interaction, and thus requires taking no IP risks from Arabic publishing houses.
We discuss TermBase creation in Section 2, the methodology in Section 3, briefly cover related work in  4, present and discuss our results in Section 5. We then discuss future directions and conclude.
@@ 2 Dataset creation
The source dataset is derived specifically from books published by our publishing house. These books span multiple academic fields, offering a diverse range of terminologies. The dataset includes terms in foreign languages (usually enclosed in parentheses) alongside their surrounding Arabic text.
The labeling process involved extracting candidate Arabic terms preceding each foreign term and annotating them with a binary label: True if the candidate is the correct translation of the foreign term, and False otherwise. To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed to correct the errors.
@@ 2.2 Annotated dataset construction
Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 5.1 to facilitate the labeling process. This draft was then reviewed to ensure the accuracy of the labels.
@@ 3.1.1 Semantic similarity
We utilized the sentence transformer LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
@@ 3.1.6 Phonetic similarity
4 Related work
Besides bilingual dictionaries, efforts have been deployed to develop lexical sources from several original languages into Arabic. Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields. Other more specialized glossaries have been developed such as the work of Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy . Furthermore, glossaries offering support to interpreters Sharif have also been established.
@@ 5 Results
Table 4 shows statistics of the Arabic books we have processed and the annotated dataset we have built to train and evaluate our work.
Table 4: Statistics from processed books
Arabic books 496
Unique foreign terms extracted 58,300
All occurrences of foreign terms 83,516
Total candidates 301,326
Average number of candidates per occurrence 3.61
Number of annotated books 9
Number of foreign terms in annotated dataset 3,595
Number of candidates in annotated dataset 15,865
@@ 5. Part of speech score $S_{POS}$:
After that, we select the candidates (of the same occurrence of a foreign term) with the highest scores within a maximum deviation (taken 0.03 here). Table 5 shows the results achieved by the heuristics approach on the whole dataset. Note that fine-tuning was done only by observing the training data. These decent results provided a good starting point that later helped in the labeling process of the data.
@@ 5.2 Machine learning
[…]
@@
As shown in Table 4, We annotated 15,865 candidates to build a training dataset. Then we trained a binary classifier using the Auto-WEKA package on the Weka Workbench Frank et al. (2016). We allowed Weka to run for several hours on 8 threads, which resulted in the following:
{strip}
Best classifier: weka.classifiers.trees.RandomForest
Arguments: [-I, 131, -K, 28, -depth, 0]
Attribute search: weka.attributeSelection.GreedyStepwise
Attribute search arguments: [-C, -R]
Attribute evaluation: weka.attributeSelection.CfsSubsetEval
Attribute evaluation arguments: [-M]
@@
Table 5 shows the results.
We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction instead of fixing a threshold, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False. This achieved a 92% precision and 98.4% recall on the total dataset, which is worse than the 99% precision and 98.5% recall achieved by the direct application of the Random Forest on the total dataset.
@@
Heuristics Precision 87%
Recall 89%
F1 Score 88%
ML Precision 100% 94.9%
@@
F1 Score 100% 93.7%
Dataset Size 12,692 (80%) 3,173 (20%)
@@ 6 Conclusion
This paper introduced methods for extracting terminology mappings between Arabic and foreign languages from Arabic books. The approach achieves useful results, contributing to automated Terminology Management and text processing. The created dataset and existing corpora further enrich the research community, advancing cross-lingual knowledge exchange.
Limitations

### v2

# MASRAD: Arabic Terminology Management Corpora with Semi-Automatic Construction

@@ Abstract
This paper presents MASRAD, a terminology dataset for Arabic terminology management, and a method with supporting tools for its semi-automatic construction. The entries in MASRAD are $(f,a)$ pairs of foreign (non-Arabic) terms $f$, appearing in specialized, academic and field-specific books next to their Arabic $a$ counterparts. MASRAD-Ex systematically extracts these pairs as a first step to construct MASRAD. MASRAD helps improving term consistency in academic translations and specialized Arabic documents, and automating cross-lingual text processing. MASRAD-Ex leverages translated terms organically occurring in Arabic books, and considers several candidate pairs for each term phrase. The candidate Arabic terms occur next to the foreign terms, and vary in length. MASRAD-Ex computes lexicographic, phonetic, morphological, and semantic similarity metrics for each candidate pair, and uses heuristic, machine learning, and machine learning with post-processing approaches to decide on the best candidate. This paper presents MASRAD after thorough expert review and makes it available to the interested research community. The best performing MASRAD-Ex approach achieved 90.5% precision and 92.4% recall.
MASRAD: Arabic Terminology Management Corpora with Semi-Automatic Construction
@@ 1 Introduction
Terminology management concerns (i) specifying terms used upon the introduction of novel concepts to a language, and (ii) enforcing the use, hence (iii) ensuring consistency and correctness. Translation from foreign languages to Arabic sources significant terminology inconsistency due to the diversity of translation experts and the lack of reference parallel term corpora. Recent systems like TURJUMAN Nagoudi et al. (2022) have improved Arabic machine translation quality, but do not focus on terminology alignment or consistency. Linguistic and structural complexities affect consistency and clarity of translations and may hinder access to knowledge.
The first and most important step is to build a termbase and provide it as a reference to Arabic scholars and experts.
We present MASRAD, a large dataset of parallel terms with pairs of foreign terms and their corresponding Arabic counterparts, sourced into use in professional, academic and scholastic Arabic contexts. We make MASRAD available for the research community and to authors looking for consistent Arabic terms when introducing a novel concept to Arabic.
@@
Recent advances in natural language processing (NLP) and computational linguistics (CL) have paved the way for automated systems that assist in cross-lingual text processing. However, existing systems, including LLMs, often struggle with the nuances of terminology in cross-lingual specific expert fields, where precision is paramount. This paper constructs MASRAD, and presents a novel methodology and a supporting framework (MASRAD-Ex) to extract and curate parallel foreign-Arabic terms from existing Arabic books.
@@
When a foreign term occurs in Arabic scholarly writing, it usually occurs between parentheses, and it is usually preceded by its Arabic translation. Two problems arise. (i) The exact scope of the Arabic term is ambiguous, and (ii) often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones. Editors and reviewers, who are usually expert researchers themselves, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
@@
MASRAD and MASRAD-Ex help automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
@@
The problem is formulated as follows. Consider a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$ between parentheses, and preceded from right to left with a sequence of Arabic words $w_{1},\ldots w_{n}$. We denote the candidate terms possible to match $f$ as the set of strings $S(f)=\{s_{i}:s_{i}=w_{1}\ldots w_{i},i\in[1,n]\}$ ending before term $f$. MASRAD-Ex extracts these candidates from the books, and identifies the target candidate term $s_{t}\in S(f)$ best matching $f$. MASRAD-Ex extracts linguistic features for each candidate match $(s_{i},f)$. It interpolates the features and aggregates their values into a resulting score, and then ranks the candidate terms.
@@
This work addresses the challenge of terminology mapping and reduces language barriers, facilitating a greater exchange of knowledge and ideas across languages and cultures.
@@
We make MASRAD and the constructed dataset of feature vectors (used for training/testing) available online for the research community upon request [^1] https://github.com/mnasser-dru/MASRAD. To the best of our knowledge, this approach to termbase building for Arabic is novel, and no comparable datasets currently exist.
@@ 1.1 Motivating example
Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme (The ethnocentrism) with different translations.
@@
MASRAD-Ex computers the following candidates per context. Candidates for Context 1:
@@
  • We build and provide MARSAD, a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
  • We build and evaluate MASRAD-Ex including three approaches: heuristic, machine learning (ML), and ML with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges.
The results are impressive and industry/deployment level, convincing for deployment. The final model runs locally, requires no LLM interaction, and thus requires no data exposure or IP risks.
We discuss MASRAD in Section 2, the methodology in Section 3, present and discuss results in Section 4, and cover related work in 5, We then discuss future directions and conclude.
@@ 2 Dataset creation
Table 1: Statistics of all processed books
Arabic books 495
Unique foreign terms 58,570
All occurrences of foreign terms 84,242
Total candidates 334,564
Average number of candidates per occurrence 3.97
Table 2: Statistics of MASRAD (annotated candidates)
Arabic books 15
Unique foreign terms 4,347
All occurrences of foreign terms 4,841
Total candidates 19,405
The source dataset is derived specifically from books published by a set of publishing houses that we keep anonymous for blind review. These books span multiple academic fields, offering a diverse range of terminologies. The dataset includes terms in foreign languages (usually enclosed in parentheses) alongside their surrounding Arabic text. Table 1 summarizes the data extracted from all books. Table 2 summarizes the data in MASRAD, our labeled and expert reviewed data that will be used later for training MASRAD-Ex.
The labeling process involved extracting candidate Arabic terms preceding each foreign term and annotating them with a binary label: True if the candidate is the correct translation of the foreign term, and False otherwise. To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed by experts to correct.
@@ 2.2 Annotated dataset construction
Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 4.1 to facilitate the labeling process. This draft was then reviewed to correct errors.
@@ 3.1.1 Semantic similarity
We utilized the sentence transformer Reimers and Gurevych (2019) LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
@@ 5. Part of speech score $S_{POS}$:
After that, we select the candidate with the maximum score within an occurrence of a foreign term. Table 7 shows the results achieved by the heuristics approach on the whole dataset. Note that fine-tuning was done only by observing the training data. These decent results provided a good starting point that later helped in the labeling process of the data.
@@ 4.2 Machine learning
$\text{x\_rank}(j)=A_{\text{occ}(j)}^{x}.\text{index\_of}(x(j))$
$\text{x\_difference}(j)=\max(A_{\text{occ}(j)}^{x})-x(j)$
@@
As shown in Table 1, We annotated 19,405 candidates to build the dataset. Then we trained a binary classifier using the Auto-WEKA package on the Weka Workbench Frank et al. (2016). We allowed Weka to run for several hours on 8 threads, which resulted in the arguments displayed in Table 6.
Table 6: AutoML arguments
Best classifier: RandomForest
Arguments: [-I, 252, -K, 0, -depth, 0]
Attribute search: GreedyStepwise
Attribute search arguments: [-C, -B, -R]
Attribute evaluation: CfsSubsetEval
Attribute evaluation arguments: [-L]
@@
Table 7 summarizes the results. The training set comprises data from nine books, and the test set comprises data from six books. We split the data at the book level to make the task more challenging, where the model is tested on books it has never seen before, which better reflects a real-world scenario.
We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False. This achieved a 90.1% precision and 92.9% recall.
@@
Heuristics Precision 88.8% 81.5%
Recall 88.2% 84.4%
F1 Score 88.5% 82.9%
ML Precision 100% 90.5%
@@
F1 Score 100% 91.4%
Dataset Size 15865 3540
@@ 4.3 Discussion and error analysis
4.4 Evaluation against a manual glossary
Over 495 books, our extraction pipeline identified 58,570 unique foreign terms. Intersecting these with our manual in‐house glossary of 46,787 concepts (a concept here consists of an English, a French, and an Arabic term) yielded 7,368 matches (15.74% of the glossary). This low overlap highlights a major problem in terminology management: assuming in the worst case that a concept was matched with both the English and French foreign terms of our extracted data, we get that at least 43,834 (74.84%) of the extracted terms are novel and need to be handled.
Focusing on the 7,368 matched concepts, we evaluated our ranking model’s ability to surface the expert Arabic term. The correct glossary entry appeared as the top automated suggestion 30.6% of the time (top-1 accuracy), within the top two automated suggestions 32.99% of the time (top-2 accuracy), and within the top three automated suggestions 33.3% of the time (top-3 accuracy).
By observing these accuracies, we can assume that the percentage of manual Arabic terms in these matched concepts that are also present in the 495 processed books to be close to our top-3 accuracy (33.3%). This reasoning is due to the minimal jumps between the top-k accuracies. From this we deduce two key insights:
  • About 66.7% of the matched concepts have a different Arabic translation by the authors or translators of the processed books than that of the expert. This greatly highlights the inconsistency problem in terminology.
  • Among the matched concepts with a present Arabic term in the books, our model placed it first in 91.89% of the cases (top-1 accuracy ÷ top-3 accuracy). This indicates that our model effectively elevates the expert translation when it is available. Note that the model performs really well when applying it to problems since it aggregates and compares results from the candidates of all occurrences of a term.
5 Related work
TURJUMAN Nagoudi et al. (2022) is a neural machine translation toolkit that translates into Modern Standard Arabic using fine-tuned AraT5 models. While effective at sentence-level translation, it does not address terminology alignment or consistency in specialized texts.
In addition to bilingual dictionaries, efforts have been made to develop lexical sources from several original languages into Arabic. Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields. Other more specialized glossaries have been developed, such as Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy . Furthermore, glossaries offering support to interpreters Sharif have also been established.
@@ 6 Conclusion
This paper introduced the semi-automatic construction of MASRAD using MASRAD-Ex, a set of tools for extracting and aligning terminology. Our work achieves useful results, contributing to automated terminology management and text processing. The created dataset and corpora further enrich the research community, advancing cross-lingual knowledge exchange.
Limitations and future work
@@
As shown in Section 4.4, 84.26% of the entries in the manual glossary were not matched. This highlights the richness of the world of terminology and the necessity of drawing on additional sources to expand our termbase. To this end, we plan to apply MASRAD-Ex to newly acquired books and expand our tooling to support other type of sources.
To improve the semantic similarity measure 3.1.1, we can fine-tune a sentence transformer Reimers and Gurevych (2019), or train a small model that augments the space of an existing one.